OpenStudy (anonymous):

simplify (1/625) ^ -3/4

OpenStudy (anonymous):

125

OpenStudy (anonymous):

how do you get that?

OpenStudy (anonymous):

(1/625) ^ -3/4 equals1/ (1/625) ^ 3/4

OpenStudy (anonymous):

without plugging it into a calculator... theres a formula we used in class for geometry

OpenStudy (anonymous):

625 = 5^4 so (5^-4)^(-3/4) = 5^(12/4) = 125